Hugh, Frank, and Luis are the only three buyers of gold in a small mining town. Their inverse demand functions for gold are aSuppose the quantity demanded of gold in this market is 3.56 ounces. Then the market price of an ounce of gold must be
Answer #1

The demand functions are, p = 256 – 329 0x = 8 -5.. (1) p= 64 -897 Q=8 - ...(2) p=16-29 Q: =8--p....(3) Compute total demandQ=8-, where 64< p < 256 1. Suppose the price of gold is $4. The total ounces of gold demanded in this market is, Q=24 - 21 (4